OpenStudy (anonymous):

9 5/8 + 4 2/3 please answer

OpenStudy (anonymous):

cthxs

myininaya (myininaya):

so we have \[9+\frac{5}{8}+4+\frac{2}{3}\] \[9+4+\frac{5}{8}+\frac{2}{3}\] \[13+\frac{5}{8}+\frac{2}{3}\] so lets look at this part of the expression: \[\frac{5}{8}+\frac{2}{3}\] how do we add fractions like these?

myininaya (myininaya):

we need to find a ______ denominator /\ | what goes in this blank?

OpenStudy (anonymous):

hmmm 5

myininaya (myininaya):

common

myininaya (myininaya):

we need to find a common denominator

OpenStudy (anonymous):

oh yeaaaa

myininaya (myininaya):

do you know what lcm means?

OpenStudy (karatechopper):

can i answer this one!?!?1

OpenStudy (anonymous):

least common denominator

OpenStudy (karatechopper):

well an lcm for this is 24

myininaya (myininaya):

ok go ahead chopper

OpenStudy (karatechopper):

yay!

myininaya (myininaya):

do you want to finish helping her or do you want me to?

OpenStudy (karatechopper):

well if u can help? i will show steps

OpenStudy (anonymous):

i want you to myiniayana its ok karate chopper let myinianyaya

OpenStudy (karatechopper):

i know thts y i said myin to help and i will show the steps

myininaya (myininaya):

\[\frac{5}{8}=\frac{?}{24}\]

myininaya (myininaya):

so lcm=24 so said the chopper

myininaya (myininaya):

so what goes in place of that question mark

OpenStudy (anonymous):

120

myininaya (myininaya):